Diego Company manufactures one product that is sold for $79 per unit in two geographic regions—the East and West regions. The14. Diego is considering eliminating the West region because an internally generated report suggests the region’s total gross margin in the first year of operations was $80,000 less than its traceable fixed selling and administrative expenses. Diego believes that if it drops the West region, the East region's sales will grow by 3% in Year 2. Using the contribution approach for analyzing segment profitability and assuming all else remains constant in Year 2, what would be the profit impact of dropping the West region in Year 2?

15. Assume the West region invests $40,000 in a new advertising campaign in Year 2 that increases its unit sales by 20%. If all else remains constant, what would be the profit impact of pursuing the advertising campaign?
Answer #1
Selling price 79
Less: Variable expense per unit 51 =29+16+2+4
Contribution margin per unit 28
14
Contribution margin lost of West division -280000 =10000*28
Contribution margin increase of East division 29400 =35000*3%*28
Avoidable fixed expenses 240000
Change in profits -10600
Profits will decrease by 10600
15
Contribution margin increase of West division 56000 =10000*20%*28
Less: Advertising costs -40000
Change in profits 16000
Profits will increase by 16000